OpenStudy (anonymous):

center and radius of a circle x^2-10x+y^2=8

OpenStudy (anonymous):

you need to "complete the square" to write it as \[(x-h)^2+(y-k)^2=r^2\] and then read off the solution as center \[(h,k)\] radius \[r\]

OpenStudy (anonymous):

(x-5)^2-25+y^2==8 (x-5)+y^2=33 (5,0) center radius sqrt(33)

OpenStudy (anonymous):

as follows \[(x-5)^2+y^2=8+5^2\] \[(x-5)^2+y^2=33\] so center is \[(5,0)\] \[r=\sqrt{33}\]

OpenStudy (anonymous):

what jahtoday said!

OpenStudy (anonymous):

yay!